Hi, I need your help on this question:
In a fish shop,60% of the fish were guppies and the rest were angelfish. After 81 fishes were removed, the number of guppies decreased by 25% and the number of angelfish was 1/4 of the original number of angelfish.
a)How many guppies were removed?
b)How many angelfish were there at first?
Thank you, I really appreciate your help!

Answer by mananth(16949)

In a fish shop,60% of the fish were guppies and the rest were angelfish. After 81 fishes were removed, the number of guppies decreased by 25% and the number of angelfish was 1/4 of the original number of angelfish.
a)How many guppies were removed?
b)How many angelfish were there at first?
Thank you, I really appreciate your help!
let total fish be x
0.6x = guppies
0.4x = angelfish
..
x-81 = 0.75*0.6x +0,25*0.4x
x-81=0.45x+0.1x
x-81=0.55x
0.45x=81
x=180 fish
60 % of 180 = 108 were guppies
40%of 180= 72 were angel fish
...
25%of 108 = 27 guppies were removed